OpenStudy (anonymous):

Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 80 miles per hour and Train B is traveling at 80 miles per hour. Train A passes a station at 10:20am if train B passed the same station at 10:50am. When will Train B catch up with A? My answer is 6:30pm

OpenStudy (anonymous):

@saifoo.khan yes you can

OpenStudy (saifoo.khan):

@pfenn1 will solve this. :D

OpenStudy (pfenn1):

If the 2 trains are traveling at the exact same speed in the same directrion, they will always stay the same distance apart, won't they?

OpenStudy (anonymous):

@saifoo.khan you can help,lol

OpenStudy (pfenn1):

sorry, @saifoo.khan, I'll go away

OpenStudy (saifoo.khan):

Should i try? i dont want to.. i hate this problem too. -_-

OpenStudy (anonymous):

@saifoo.khan but your a math wiz,lol

OpenStudy (anonymous):

So, my answer is wrong,huh?

OpenStudy (saifoo.khan):

http://media.tumblr.com/tumblr_lxcx34cYlH1r54pcz.jpg

OpenStudy (saifoo.khan):

let me solve.

OpenStudy (anonymous):

Pic is tooo funny

OpenStudy (saifoo.khan):

Lol. ;) so are both terms 80 mph and 80mph?

OpenStudy (anonymous):

B is 90 miles per hour

OpenStudy (anonymous):

Sorry about that

OpenStudy (saifoo.khan):

No prob.

OpenStudy (saifoo.khan):

this i what im getting. im not sure. :( 10:20 + 3 = 13:20 or in other words, 1:20pm.

OpenStudy (anonymous):

So, my 6:30pm is wrong,lol

OpenStudy (anonymous):

I'm not sure what the 13:20 means

OpenStudy (saifoo.khan):

1320 hrs (in 24 hours clock)

OpenStudy (anonymous):

ha oh, but where are you getting the 3?

OpenStudy (anonymous):

it's 10 difference right?

OpenStudy (saifoo.khan):

How i got 3? Answer = 30/60 * 90 = 45 45/10 = 4.5hrs = 4 hrs and 30mins therefore, 10:20 + 4:30 02:50 pm.

OpenStudy (saifoo.khan):

Sorry i did a mistake.

OpenStudy (anonymous):

Wow, I just checked my work again and got 3:50pm, Let me check again

OpenStudy (saifoo.khan):

Sure.

OpenStudy (anonymous):

I'm still getting a wrong answer but since your the math wiz @saifoo.khan I am going with your answer if 2:50pm lol

OpenStudy (saifoo.khan):

P.S. my answer might be wrong. i always skip this question! ;D

OpenStudy (anonymous):

hahaahaha....hey so my answer of 6:30pm may still be right,lol

OpenStudy (saifoo.khan):

Haha, there can be chances.. ain't gonna say anything.

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

One more question then I'm off to bed...math is going to be the death of me,lol

OpenStudy (saifoo.khan):

Haha, there are more weeks to come! you sure will enjoy. ;D